A grocery store manager would like to have an idea concerning the average amount milk the store sells per day. In a sample of 77 days, the average amount number of gallons sold was 125 with a standard deviation of 24 gallons. What is the upper bound of the 95% confidence interval? (please express your answer using 2 decimal places)
Answer #1

CI = Confidence interval

CI upper bound = X + z (SD/ n​​​​​​0.5)

SD = Standard deviation = 24

X = average amount = 125

z = 1.96 for 95% CI level

n = 77 days

95% CI upper bound = 125 + (1.96)(24/(77)0.5)

95% CI upper bound = 125 + (1.96)(2.74) = 125+5.37

95% CI upper bound = 130.37